### v3

@@ 3.2.1 Subspace-Based Adaptation
Subspace-based domain adaptation (DA) methods typically project source and target samples into separate subspaces using subspace learning or dimensionality reduction techniques, and then align the subspaces. The Subspace Alignment (SA) method proposed by Fernando et al. [28] has been extended to multispectral agricultural image classification tasks. By employing Principal Component Analysis (PCA), this method aligns crop spectral feature subspaces across different seasons, thereby alleviating domain shifts caused by lighting and phenological variations. The SA process is illustrated in Figure 5.
@@
As shown, the source domain samples $X_{S}\in\mathbb{R}^{n_{s}\times D}$ contain $n_{s}$ samples with $D$-dimensional features, and the target domain samples $X_{T}\in\mathbb{R}^{n_{t}\times D}$ share the same dimensionality. PCA is applied to obtain the principal components of source and target domains, and a linear transformation matrix $M$ is learned for subspace alignment.
@@
[…]
where $P_{s}$ and $P_{t}$ are the projection matrices (e.g., PCA bases) of the source and target subspaces, respectively, and $\|\cdot\|_{F}^{2}$ denotes the Frobenius norm used to measure domain discrepancy.
Sun et al. [29] applied SA to cross-view scene classification using Partial Least Squares (PLS) to construct a discriminative subspace shared by both domains, effectively capturing common spectral features across different sensors. They further proposed the Transfer Sparse Subspace Analysis (TSSA) method [30], which incorporates sparse subspace clustering and Maximum Mean Discrepancy (MMD) minimization to retain self-expressive properties of data across maize growth stages.
Weilandt et al. [31] introduced the Early Crop Mapping based on Dynamic Clustering Method (ECMDCM), leveraging time-series NDVI and EVI data for improved accuracy in early-stage crop classification. Li et al. [32] proposed CropSTGAN, which includes a domain mapper to adapt spatial-temporal discrepancies in crop mapping, while Wang et al. [33] developed MultiCropGAN, incorporating identity loss to preserve essential features during label space adaptation, with validation in North American farmlands.
Adversarial learning techniques based on category alignment offer another avenue. Takahashi et al. [34] proposed a category-guided feature alignment framework, which maps features from both domains into a shared semantic space using adversarial training. Without requiring target domain labels, it leverages source domain categories to enhance domain-invariant learning, improving generalization across crop types and acquisition conditions.
@@
To address challenges in hyperspectral image classification, such as label-feature inconsistency and computational inefficiency in similarity learning, Yi [39] proposed a subspace learning framework based on set similarity metrics. Compared with traditional methods like Linear Discriminant Analysis (LDA), this approach improves classification accuracy by 8%–15% and reduces training time by 40%–60% on datasets such as Indian Pines and Salinas Valley.
Finally, Banerjee [40] proposed a hierarchical subspace alignment method using a semantic-driven binary tree and Grassmannian manifold alignment to address multi-temporal distribution mismatches due to atmospheric, sensor, and labeling differences. It outperforms traditional techniques (e.g., Global GFK, SA) by 5%–6% in accuracy and substantially reduces misclassification in semantically similar crop categories, offering a scalable solution for dynamic agricultural monitoring and environmental analysis.
@@ 3.2.2 Transformation-Based Adaptation
Aptoula et al. [41] proposed a deep learning–based framework for domain-adaptive classification of crops and weeds, addressing distributional differences caused by variations in lighting, soil properties, and plant phenology. This method embodies the essence of transformation-based DA by projecting agricultural remote sensing data into a new feature space that minimizes both statistical and geometric distributional differences across domains, while preserving structural consistency. These methods have demonstrated substantial utility in agricultural remote sensing and crop monitoring applications (see Figure 6).
@@
The transformation process typically relies on metrics such as Maximum Mean Discrepancy (MMD), Kullback-Leibler (KL) divergence, or Bregman divergence to quantify inter-domain differences. Among these, MMD is widely favored for agricultural data due to its robustness in high-dimensional settings, and is defined as:
@@
  • Polynomial Kernel: Can model nonlinear feature relationships, especially when inter-band correlations exist; however, high-degree polynomials may incur overfitting and computational costs.
@@
In general, the RBF kernel remains the preferred choice for hyperspectral and other complex agricultural data types due to its robust performance in capturing nonlinear structures.
Pan et al. [42] introduced Transfer Component Analysis (TCA), which projects source and target data into a shared RKHS and minimizes domain discrepancy via MMD. The TCA objective function is:
@@ 3.3 Classifier-Based Adaptation
In agriculture, domain shifts often arise due to seasonal changes, varying crop phenotypes, and different imaging conditions. To address these, Bruzzone et al.[43, 44, 45, 46] introduced a maximum likelihood classifier adaptation method, which iteratively updates classifier parameters to better reflect the statistical properties of new-season imagery. Similarly, an improved Bayesian cascade classifier[44] enhanced the robustness of cross-regional crop-type recognition. Zhong et al.[47] integrated spectral-spatial guided filtering with adaptive classifier updating, significantly boosting the detection accuracy of crop diseases in heterogeneous field conditions.
Given the complexity of cross-domain agricultural data, ensemble learning has gained traction as an effective strategy to increase robustness. By aggregating predictions from multiple base classifiers, ensemble techniques mitigate overfitting and model bias. Wei et al.[48] proposed a Multi-Domain Adaptation Fusion (MDAF) method that improves farmland classification accuracy across different countries through classifier ensembles, addressing spectral variability.
Weighted classifier transfer approaches—especially those based on Random Forests—are also extensively employed in agriculture. Zhang et al.[49] utilized weighted Random Forests to merge agricultural data from distinct geographic regions, enhancing the performance of cross-regional crop disease detection. Xu et al.[50] extended Extreme Learning Machine (ELM) transfer with kernel alignment to predict crop yields under varying climate conditions, further exemplifying the utility of classifier-based adaptation in agriculture.
To cope with the high cost of labeling agricultural imagery, semi-supervised learning (SSL) and active learning (AL) have become increasingly prominent. Rajan et al.[51] proposed a Binary Hierarchical Classifier (BHC) integrating SSL for continuous model updates in cross-year pest monitoring. Additionally, enhancements to the classic Domain Adaptation SVM (DASVM) have enabled improved generalization in cross-sensor crop classification with minimal labeled target samples[52].
@@
Figure 7 illustrates a typical AL process. Initially, a classifier is trained on the source domain and used to label the target domain. Samples with the highest prediction uncertainty are selected for manual labeling and added back to the training set, refining the classifier iteratively. AL is highly effective in precision agriculture due to its ability to reduce annotation demands.
Deng et al.[53] developed an Active Multi-Kernel Domain Adaptation (AMKDA) method combining AL with multi-kernel learning, substantially reducing the need for labeled hyperspectral data. Kalita et al.[54] introduced a standard deviation (SD)-based AL strategy that identifies the most informative samples, while Saboori et al.[55] proposed Active Multi-Kernel Fredholm Learning (AMKFL), which uses Fredholm regularization to guide annotation and improve classification.
@@ 4.1 Supervised Deep DA
Supervised deep DA methods assume access to a limited amount of labeled data in the target domain. These methods jointly optimize the feature extractor and classifier by minimizing the combination of source domain classification loss and cross-domain alignment loss. The primary idea is to construct a deep neural network that integrates target domain label supervision into the domain alignment process. A typical supervised deep DA model consists of three modules: a feature extractor, a classifier, and a domain discrepancy measurement component. The joint loss function is formulated as:
@@
Here, $L_{cls}(X_{s},Y_{s})$ denotes the classification loss on the source domain, $L_{align}(X_{s},X_{t})$ measures the domain discrepancy (e.g., via MMD, adversarial loss, or covariance alignment), and $\lambda$ is a trade-off parameter.
In agricultural scenarios, supervised deep DA is widely used for cross-sensor data fusion and multi-temporal crop monitoring. For instance, Takahashi et al. [34] proposed a category-aware adversarial learning framework that utilizes source domain class labels to guide feature space alignment, thereby improving generalization on target agricultural datasets collected from heterogeneous imaging devices (see the Figure 8).
@@
This approach comprises three main stages: image generation, transfer learning, and domain adaptation based on deep neural networks (DNNs). First, a Class Propagation Network (CPN) generates a class map from target images, simulating realistic agricultural image distributions. Second, transfer learning is performed on both real images ($I$) and synthetic images ($W$), producing two transfer models, TL-I and TL-W. Finally, deep networks implement domain adaptation using two modules: DA-W2I (synthetic-to-real adaptation) and DA-I2W (real-to-synthetic adaptation).
As shown in Figure 8, the system architecture includes five components: CPN for image generation, TL-I and TL-W for transfer learning, and DA-W2I and DA-I2W for bi-directional domain adaptation [58]. These modules synergistically enhance generalization on the agricultural target domain, especially under limited label availability, leveraging both image-level and feature-level adaptation strategies.
Similarly, Espejo-García et al. [59] proposed a method that integrates agricultural transfer learning with Generative Adversarial Networks (GANs) to synthesize annotated images, addressing the scarcity of labeled data in agricultural datasets.
@@ 4.2 Semi-supervised Deep Domain Adaptation (DA)
Self-training and pseudo-label generation are core techniques in semi-supervised DA. Recent innovations have integrated diffusion models into this paradigm, showing promising results in agricultural applications. Ghanbari et al. [65] proposed a framework that combines diffusion models with meta-learning to perform wheat ear segmentation using only a few annotated samples. Specifically, the method requires just three manually labeled wheat images and unlabelled video frames to generate a large pseudo-labeled dataset through a probabilistic diffusion process.
The architecture features a dual-branch encoder-decoder network designed to enable effective feature alignment and semantic consistency between domains (see Figure 9). The model achieved a Dice score of 63.5% on an external test set comprising farmland images from multiple countries, demonstrating strong generalization capabilities. This diffusion-based pseudo-labeling mechanism substantially improves the adaptability of deep DA in data-scarce agricultural environments, offering both scalability and practical relevance.
@@
In this architecture, the encoder learns joint representations for both synthetic and real images, while the mask decoder generates segmentation masks. The image decoder reconstructs real images, encouraging the encoder to adapt to real domain features. Experimental results demonstrate Dice scores of 80.7% on internal test sets and 64.8% on external test sets from 18 plots across five countries. These results highlight the framework’s potential to support generalizable agricultural vision models and facilitate real-world deployment.
By incorporating diffusion models, this approach effectively harnesses large-scale unlabeled datasets and has shown strong performance in image processing tasks [66, 67], underscoring its relevance in precision agriculture.
To address wheat disease detection, Najafian et al. [68] developed the WheatSeedBelt dataset and designed a pipeline combining pretrained models with semi-supervised fine-tuning, achieving high accuracy in Fusarium-damaged kernel (FDK) classification. Subsequently, they proposed a semi-self-supervised wheat head detection method [69], using few annotated images and video sequences to generate pseudo-labeled datasets enhanced by domain adaptation techniques (mAP = 0.827).
@@
In vegetable detection, Yang et al. [71] proposed TDA-YOLO, a YOLOv5-based model adapted for tomato detection in densely planted fields under variable lighting. The model employs neural preset color style transfer to generate a pseudo dataset, narrowing the domain gap. Semi-supervised learning and knowledge distillation further enhance target domain adaptability, while the lightweight FasterNet backbone improves inference speed.
Zhu et al. [72] introduced Semi-supervised Center-based Discriminative Adversarial Learning (SCDAL) for domain-adaptive aerial scene classification. Center loss-guided adversarial learning strengthens domain discriminability.
Teng et al. [73] proposed the Classifier-constrained Deep Adversarial Domain Adaptation (CDADA) method, which integrates the Maximum Classifier Discrepancy (MCD) framework with deep convolutional neural networks (DCNNs) to enable semi-supervised cross-domain classification in remote sensing by adversarially aligning feature distributions.
@@ 4.3.1 Adversarial Training
Adversarial learning, particularly via Generative Adversarial Networks (GANs), is widely employed in Unsupervised Domain Adaptation (UDA), primarily for aligning the data distributions of the source and target domains. Two of the most influential adversarial training techniques are the Domain-Adversarial Neural Network (DANN) and the Gradient Reversal Layer (GRL), both of which are conceptually derived from the GAN framework.
@@
A GAN consists of two neural networks in competition [83]:
@@
$\min_{G}\max_{D}V(D,G)$ $=\mathbb{E}_{x\sim P_{\mathrm{data}}(x)}[\log D(x)]$
$\quad+\mathbb{E}_{z\sim P_{z}(z)}[\log(1-D(G(z)))]$ (6)
@@
A typical GAN consists of two components: a generator and a discriminator. The generator is trained to produce target domain samples based on the distribution of source domain data, while the discriminator attempts to distinguish real samples from generated ones. In the agricultural domain, GANs have been extensively used for cross-sensor image style transfer. For instance, CycleGAN has been applied to transform the visual style of source domain remote sensing imagery to resemble that of the target sensor, thus mitigating domain shifts [84].
Zhang [85] proposed an unsupervised adversarial domain adaptation method for agricultural land extraction, introducing a multi-scale feature fusion (MSFF) module to adapt across datasets with varying spatial resolutions. This design enhances the model’s robustness in learning domain-invariant features.
@@
Ji et al. [86] introduced an end-to-end GAN-based DA framework for land cover classification using multi-source imagery. In this approach, source images are translated into the target style via adversarial learning, and semantic segmentation is performed using a fully convolutional network (FCN). Similarly, Tasar et al. [87] proposed a standard GAN-based DA pipeline for semantic segmentation of very high-resolution (VHR) satellite images and extended it to a multi-source, multi-target, and lifelong learning setting. Wittich et al. [88] designed a Color Map GAN for consistent appearance adaptation in aerial image classification. Liu et al. [89] developed a GAN-based feature extractor for agricultural scene classification that enhances cross-domain classification accuracy by aligning source and target distributions.
While GANs perform exceptionally well in unpaired style transfer tasks, their major drawbacks include susceptibility to mode collapse and high computational cost during training.
On the other hand, Discriminative Adversarial Networks, such as DANN, incorporate a domain classifier that distinguishes between source and target domain features. During training, a gradient reversal layer enables backpropagation to align feature distributions implicitly, thus improving the generalization capacity of the learned model in the target domain. In agriculture, DANN has been widely adopted for domain-invariant feature extraction and alignment.
@@
Additional applications of DANN in agricultural image analysis include Bejiga et al. [92], who applied a DANN-based model for large-scale land cover classification using multispectral data. Rahhal et al. [93] designed a multi-source DA strategy using a min-max entropy optimization technique. Elshamli et al. [94] combined denoising autoencoders (DAEs) with DANN to handle DA in multi-temporal and multi-spatial imagery. Mauro et al. [95] integrated self-attention mechanisms into adversarial networks to address domain variance in multi-temporal land cover classification. Mateo et al. [96] proposed a cross-sensor adversarial DA method for cloud detection using Landsat-8 and Proba-V imagery. Ma et al. [97] introduced an Adaptive DANN (ADANN) framework that aligns meteorological features and vegetation indices across U.S. agricultural zones, achieving a 34.1% reduction in yield prediction RMSE.
Discriminative adversarial networks offer the advantage of low annotation dependency while significantly enhancing cross-domain generalization. However, they are sensitive to feature extractor quality and may exhibit convergence instability under large domain gaps.
Adversarial DA strategies, including hybrid GAN-DANN frameworks, are being widely used in remote sensing tasks such as scene classification [84], crop classification [98], road extraction [83], and cloud detection [99]. Nonetheless, the use of fixed distance metrics (e.g., MMD) may not suffice for diverse agricultural imagery datasets.
Recent advances further refined adversarial DA. Hoffman et al. [100] first integrated adversarial training into pixel-level domain adaptation. Tsai et al. [101] proposed AdaptSegNet to reduce domain shifts at the segmentation output level. Vu et al. [102] introduced ADVENT, which applies entropy-based adversarial loss to segmentation outputs. Li et al. [103] proposed a bi-directional learning (BDL) approach to enhance bidirectional feature alignment. Zhang et al. [104] designed DFENet, a domain feature enhancement network that models both channel and spatial dependencies to improve discriminability. Zheng et al. [105] presented TSAN, a two-stage adaptive network for multi-target remote sensing scene classification. Rahhal et al. [93] again applied multi-source adversarial adaptation using entropy-based alignment. Makkar et al. [106] used adversarial learning to extract target-discriminative geospatial features.
Finally, Saito et al. [107] proposed an open-set adversarial DA framework that explicitly separates known and unknown target classes. During training, the feature generator either aligns the sample with known source categories or rejects it as an unknown sample, effectively supporting open-set classification in domain shift scenarios.

### v4

@@ 3.2.1 Subspace-Based Adaptation
Subspace-based domain adaptation (DA) methods typically project source and target samples into separate subspaces using subspace learning or dimensionality reduction techniques, and then align the subspaces. The Subspace Alignment (SA) method proposed by Fernando et al. [29] has been extended to multispectral agricultural image classification tasks. By employing Principal Component Analysis (PCA), this method aligns crop spectral feature subspaces across different seasons, alleviating domain shifts caused by lighting and phenological variations. The SA process is illustrated in Figure 5.
@@
As shown, the source domain samples $X_{S}\in\mathbb{R}^{n_{s}\times D}$ contain $n_{s}$ samples with $D$-dimensional features, and the target domain samples $X_{T}\in\mathbb{R}^{n_{t}\times D}$ share the same dimensionality. PCA is applied to obtain the principal components of the source and target domains, and a linear transformation matrix $M$ is learned for subspace alignment.
@@
$M=\underset{M}{\operatorname*{arg\,min}}\parallel P_{s}M-P_{t}\parallel_{F}^{2}=P_{s}^{T}P_{t}$ (2)
where $P_{s}$ and $P_{t}$ are the projection matrices (e.g., PCA bases) of the source and target subspaces, respectively, and $\|\cdot\|_{F}^{2}$ denotes the Frobenius norm.
Sun et al. [30] applied SA to cross-view scene classification using Partial Least Squares (PLS) to construct a discriminative subspace shared by both domains, effectively capturing common spectral features across different sensors. They further proposed the Transfer Sparse Subspace Analysis (TSSA) method [31], which incorporates sparse subspace clustering and Maximum Mean Discrepancy (MMD) minimization to retain self-expressive properties of data across maize growth stages. Weilandt et al. [32] introduced the Early Crop Mapping based on Dynamic Clustering Method (ECMDCM), leveraging time-series NDVI and EVI data for improved early-stage crop classification. Li et al. [33] proposed CropSTGAN, which includes a domain mapper to adapt spatial-temporal discrepancies in crop mapping, while Wang et al. [34] developed MultiCropGAN, incorporating identity loss to preserve essential features during label space adaptation, with validation in North American farmlands.
Adversarial learning techniques based on category alignment offer another promising avenue. Takahashi et al. [35] proposed a category-guided feature alignment framework, which maps features from both domains into a shared semantic space using adversarial training. Without requiring target domain labels, it leverages source domain categories to enhance domain-invariant learning, thereby improving generalization across crop types and acquisition conditions.
@@
To address challenges in hyperspectral image classification, such as label-feature inconsistency and computational inefficiency in similarity learning, Yi [40] proposed a subspace learning framework based on set similarity metrics. Compared with traditional methods like Linear Discriminant Analysis (LDA), this approach improves classification accuracy by 8%–15% and reduces training time by 40%–60% on datasets such as Indian Pines and Salinas Valley. Finally, Banerjee [41] proposed a hierarchical subspace alignment method using a semantic-driven binary tree and Grassmannian manifold alignment to address multi-temporal distribution mismatches due to atmospheric, sensor, and labeling differences. It outperforms traditional techniques (e.g., Global GFK, SA) by 5%–6% in accuracy and substantially reduces misclassification in semantically similar crop categories, offering a scalable solution for dynamic agricultural monitoring and environmental analysis.
@@ 3.2.2 Transformation-Based Adaptation
Aptoula et al. [42] proposed a deep learning–based framework for domain-adaptive classification of crops and weeds, addressing distributional differences caused by variations in lighting, soil properties, and plant phenology. This method embodies the essence of transformation-based DA by projecting agricultural remote sensing data into a new feature space that minimizes both statistical and geometric differences across domains while preserving structural consistency. As illustrated in Figure 6, these methods have shown substantial utility in agricultural remote sensing and crop monitoring applications.
@@
The transformation process typically relies on metrics such as Maximum Mean Discrepancy (MMD), Kullback-Leibler (KL) divergence, or Bregman divergence to quantify inter-domain differences. Among these, MMD is widely favored for agricultural data due to its robustness in high-dimensional settings and is defined as:
@@
  • Polynomial Kernel: Can model nonlinear feature relationships, especially when inter-band correlations exist; however, high-degree polynomials may incur overfitting and high computational costs.
@@
In general, the RBF kernel remains the preferred choice for hyperspectral and other complex agricultural data types due to its robust performance in capturing nonlinear structures. Pan et al. [43] introduced Transfer Component Analysis (TCA), which projects source and target data into a shared RKHS and minimizes domain discrepancy via MMD. The TCA objective function is:
@@ 3.3 Classifier-Based Adaptation
In agriculture, domain shifts often arise due to seasonal changes, varying crop phenotypes, and different imaging conditions. To address these challenges, Bruzzone et al. [44, 45, 46, 47] introduced a maximum likelihood classifier adaptation method that iteratively updates classifier parameters to better reflect the statistical properties of new-season imagery. Similarly, an improved Bayesian cascade classifier [45] enhanced the robustness of cross-regional crop-type recognition. Zhong et al. [48] integrated spectral-spatial guided filtering with adaptive classifier updating, significantly boosting the detection accuracy of crop diseases under heterogeneous field conditions.
Given the complexity of cross-domain agricultural data, ensemble learning has gained traction as an effective strategy to increase robustness. By aggregating predictions from multiple base classifiers, ensemble techniques mitigate overfitting and model bias. Wei et al. [49] proposed a Multi-Domain Adaptation Fusion (MDAF) method that improves farmland classification accuracy across different countries through classifier ensembles, effectively addressing spectral variability.
Weighted classifier transfer approaches—especially those based on Random Forests—are also extensively employed in agriculture. Zhang et al. [50] utilized weighted Random Forests to merge agricultural data from distinct geographic regions, enhancing the performance of cross-regional crop disease detection. Xu et al. [51] extended the Extreme Learning Machine (ELM) framework with kernel alignment to predict crop yields under varying climate conditions, further demonstrating the utility of classifier-based adaptation in agriculture.
To cope with the high cost of labeling agricultural imagery, semi-supervised learning (SSL) and active learning (AL) have become increasingly prominent. Rajan et al. [52] proposed a Binary Hierarchical Classifier (BHC) that integrates SSL for continuous model updates in cross-year pest monitoring. Additionally, enhancements to the classic Domain Adaptation SVM (DASVM) have enabled improved generalization in cross-sensor crop classification with minimal labeled target samples [53].
Figure 7 illustrates the flowchart of a typical AL method for domain adaptation. Initially, a classifier is trained on the source domain and used to label the target domain. Samples with the highest prediction uncertainty are selected for manual labeling and added back to the training set, iteratively refining the classifier. AL is highly effective in precision agriculture due to its ability to reduce annotation demands. Deng et al. [54] developed an Active Multi-Kernel Domain Adaptation (AMKDA) method combining AL with multi-kernel learning, substantially reducing the need for labeled hyperspectral data. Kalita et al. [55] introduced a standard deviation (SD)-based AL strategy to identify the most informative samples, while Saboori et al. [56] proposed Active Multi-Kernel Fredholm Learning (AMKFL), which uses Fredholm regularization to guide annotation and improve classification.
@@ 4.1 Supervised Deep DA
Supervised deep DA methods assume access to a limited amount of labeled data in the target domain. These methods jointly optimize the feature extractor and classifier by minimizing a combination of source domain classification loss and cross-domain alignment loss. The primary idea is to construct a deep neural network that incorporates target domain label supervision into the domain alignment process. A typical supervised deep DA model consists of three modules: a feature extractor, a classifier, and a domain discrepancy measurement component. The joint loss function is formulated as:
@@
Here, $L_{cls}(X_{s},Y_{s})$ denotes the classification loss on the source domain, $L_{align}(X_{s},X_{t})$ measures domain discrepancy (e.g., via MMD, adversarial loss, or covariance alignment), and $\lambda$ is a trade-off parameter.
In agricultural scenarios, supervised deep DA is widely used for cross-sensor data fusion and multi-temporal crop monitoring. For instance, Takahashi et al. [35] proposed a category-aware adversarial learning framework that leverages source domain class labels to guide feature space alignment, thereby enhancing generalization on target agricultural datasets collected from heterogeneous imaging devices.
@@
This approach comprises three main stages: image generation, transfer learning, and domain adaptation based on deep neural networks (DNNs). First, a Class Propagation Network (CPN) generates class maps from target images, simulating realistic agricultural image distributions. Second, transfer learning is performed on both real images ($I$) and synthetic images ($W$), producing two transfer models, TL-I and TL-W. Finally, deep networks implement domain adaptation using two modules: DA-W2I (synthetic-to-real adaptation) and DA-I2W (real-to-synthetic adaptation).
As shown in Figure 8, the system architecture includes five components: CPN for image generation, TL-I and TL-W for transfer learning, and DA-W2I and DA-I2W for bi-directional domain adaptation [59]. These modules synergistically enhance generalization on the agricultural target domain, especially under limited label availability, leveraging both image-level and feature-level adaptation strategies. Similarly, Espejo-García et al. [60] proposed a method that integrates agricultural transfer learning with Generative Adversarial Networks (GANs) to synthesize annotated images, addressing the scarcity of labeled data in agricultural datasets.
@@ 4.2 Semi-supervised Deep Domain Adaptation (DA)
Self-training and pseudo-label generation are core techniques in semi-supervised DA. Recent innovations have integrated diffusion models into this paradigm, showing promising results in agricultural applications. Ghanbari et al. [66] proposed a framework that combines diffusion models with meta-learning to perform wheat ear segmentation using only a few annotated samples. Specifically, the method requires just three manually labeled wheat images and unlabeled video frames to generate a large pseudo-labeled dataset through a probabilistic diffusion process.
The architecture incorporates a dual-branch encoder-decoder network designed to facilitate effective feature alignment and maintain semantic consistency between domains, as shown in Figure 9. The model achieved a Dice score of 63.5% on an external test set comprising farmland images from multiple countries, demonstrating strong generalization capabilities. This diffusion-based pseudo-labeling mechanism substantially improves the adaptability of deep DA in data-scarce agricultural environments, offering both scalability and practical relevance.
@@
In this architecture, the encoder learns joint representations for both synthetic and real images, while the mask decoder generates segmentation masks. The image decoder reconstructs real images, encouraging the encoder to adapt to real domain features. Experimental results demonstrate Dice scores of 80.7% on internal test sets and 64.8% on external test sets from 18 plots across five countries. These results highlight the framework’s potential to support generalizable agricultural vision models and facilitate real-world deployment. By incorporating diffusion models, this approach effectively harnesses large-scale unlabeled datasets and has shown strong performance in image processing tasks [67, 68], underscoring its relevance in precision agriculture.
To address wheat disease detection, Najafian et al. [69] developed the WheatSeedBelt dataset and designed a pipeline combining pretrained models with semi-supervised fine-tuning, achieving high accuracy in Fusarium-damaged kernel (FDK) classification. Subsequently, they proposed a semi-self-supervised wheat head detection method [70], using a few annotated images and video sequences to generate pseudo-labeled datasets enhanced by domain adaptation techniques (mAP = 0.827).
@@
In vegetable detection, Yang et al. [72] proposed TDA-YOLO, a YOLOv5-based model adapted for tomato detection in densely planted fields under variable lighting. The model employs neural preset color style transfer to generate a pseudo dataset, narrowing the domain gap. Semi-supervised learning and knowledge distillation further enhance target domain adaptability, while the lightweight FasterNet backbone improves inference speed. Zhu et al. [73] introduced Semi-supervised Center-based Discriminative Adversarial Learning (SCDAL) for domain-adaptive aerial scene classification. Center-loss-guided adversarial learning strengthens domain discriminability. Teng et al. [74] proposed the Classifier-constrained Deep Adversarial Domain Adaptation (CDADA) method, which integrates the Maximum Classifier Discrepancy (MCD) framework with deep convolutional neural networks (DCNNs) to enable semi-supervised cross-domain classification in remote sensing by adversarially aligning feature distributions.
@@ 4.3.1 Adversarial Training
Adversarial training, particularly via Generative Adversarial Networks (GANs), is widely employed in Unsupervised Domain Adaptation (UDA), primarily for aligning the data distributions of the source and target domains. Two of the most influential adversarial training techniques are the Domain-Adversarial Neural Network (DANN) and the Gradient Reversal Layer (GRL), both of which are conceptually derived from the GAN framework.
@@
As illustrated in Figure 10, a GAN consists of two neural networks in competition [84]:
@@
$\min_{G}\max_{D}V(D,G)=\mathbb{E}_{x\sim P_{\mathrm{data}}(x)}[\log D(x)]+\mathbb{E}_{z\sim P_{z}(z)}[\log(1-D(G(z)))]$ (6)
@@
A typical GAN consists of two components: a generator and a discriminator. The generator is trained to produce target domain samples based on the distribution of source domain data, while the discriminator attempts to distinguish real samples from generated ones. In the agricultural domain, GANs have been extensively used for cross-sensor image style transfer. For instance, CycleGAN has been applied to transform the visual style of source domain remote sensing imagery to resemble that of the target sensor, thus mitigating domain shifts [85]. Zhang [86] proposed an unsupervised adversarial domain adaptation method for agricultural land extraction, as illustrated in Figure 11, introducing a multi-scale feature fusion (MSFF) module to adapt across datasets with varying spatial resolutions. This design enhances the model’s robustness in learning domain-invariant features.
@@
Ji et al. [87] introduced an end-to-end GAN-based DA framework for land cover classification using multi-source imagery. In this approach, source images are translated into the target style via adversarial learning, and semantic segmentation is performed using a fully convolutional network (FCN). Similarly, Tasar et al. [88] proposed a standard GAN-based DA pipeline for semantic segmentation of very high-resolution (VHR) satellite images and extended it to a multi-source, multi-target, and lifelong learning setting. Wittich et al. [89] designed a Color Map GAN for consistent appearance adaptation in aerial image classification. Liu et al. [90] developed a GAN-based feature extractor for agricultural scene classification that enhances cross-domain classification accuracy by aligning source and target distributions. While GANs perform exceptionally well in unpaired style transfer tasks, their major drawbacks include susceptibility to mode collapse and high computational cost during training.
On the other hand, Discriminative Adversarial Networks, such as DANN, as illustrated in Figure 12, incorporate a domain classifier that distinguishes between source and target domain features. During training, a gradient reversal layer enables backpropagation to align feature distributions implicitly, thus improving the generalization capacity of the learned model in the target domain. In agriculture, DANN has been widely adopted for domain-invariant feature extraction and alignment.
@@
Additional applications of DANN in agricultural image analysis include Bejiga et al. [93], who applied a DANN-based model for large-scale land cover classification using multispectral data. Rahhal et al. [94] designed a multi-source DA strategy using a min-max entropy optimization technique. Elshamli et al. [95] combined denoising autoencoders (DAEs) with DANN to handle DA in multi-temporal and multi-spatial imagery. Mauro et al. [96] integrated self-attention mechanisms into adversarial networks to address domain variance in multi-temporal land cover classification. Mateo et al. [97] proposed a cross-sensor adversarial DA method for cloud detection using Landsat-8 and Proba-V imagery. Ma et al. [98] introduced an Adaptive DANN (ADANN) framework that aligns meteorological features and vegetation indices across U.S. agricultural zones, achieving a 34.1% reduction in yield prediction RMSE. Discriminative adversarial networks offer the advantage of low annotation dependency while significantly enhancing cross-domain generalization. However, they are sensitive to feature extractor quality and may exhibit convergence instability under large domain gaps.
Adversarial DA strategies, including hybrid GAN-DANN frameworks, are being widely used in remote sensing tasks such as scene classification [85], crop classification [99], road extraction [84], and cloud detection [100]. Nonetheless, the use of fixed distance metrics (e.g., MMD) may not suffice for diverse agricultural imagery datasets. Recent advances further refined adversarial DA. Hoffman et al. [101] first integrated adversarial training into pixel-level domain adaptation. Tsai et al. [102] proposed AdaptSegNet to reduce domain shifts at the segmentation output level. Vu et al. [103] introduced ADVENT, which applies entropy-based adversarial loss to segmentation outputs. Li et al. [104] proposed a bi-directional learning (BDL) approach to enhance bidirectional feature alignment. Zhang et al. [105] designed DFENet, a domain feature enhancement network that models both channel and spatial dependencies to improve discriminability. Zheng et al. [106] presented TSAN, a two-stage adaptive network for multi-target remote sensing scene classification. Rahhal et al. [94] again applied multi-source adversarial adaptation using entropy-based alignment. Makkar et al. [107] used adversarial learning to extract target-discriminative geospatial features. Finally, Saito et al. [108] proposed an open-set adversarial DA framework that explicitly separates known and unknown target classes. During training, the feature generator either aligns the sample with known source categories or rejects it as an unknown sample, effectively supporting open-set classification in domain shift scenarios.